OpenStudy (anonymous):

PLEASE HELP!! 1. A soccer player kicks the soccer ball straight up with a velocity of 17 m/s. Using a distance versus time graph: How long is the ball in the air? Using a velocity versus displacement graph: How high does it rise? 2. A dart is dropped from a height of 4.12 meters. Using a distance versus time graph How long is the dart in the air?

OpenStudy (aum):

Are you expected to graph them or is there a diagram associated with this question?

OpenStudy (anonymous):

It's a graph

OpenStudy (aum):

Can you post the graphs? Post a screenshot of the question.

OpenStudy (anonymous):

But it doesn't even show the graph... It just says Using a distance verses time graph: (whole sentence underline) and that's it

OpenStudy (anonymous):

I'm guess I'm supposed to do equations, but I'm so confused

OpenStudy (aum):

The formula is: \(y = ut - \frac 12 gt^2 = 17t - \frac 12 * 10 * t^2 = -5t^2 + 17t\) Graph y vs t. The x-intercept of the graph, other than t = 0, will be how long the ball stays in the air.

OpenStudy (anonymous):

Oh that makes a bit more sense now. I just need the equation

OpenStudy (anonymous):

Could you also explain the rest, please?

OpenStudy (aum):

Have you done calculus yet?

OpenStudy (anonymous):

I'm currently in Trig and Analysis of Functions...

OpenStudy (aum):

Have you covered the equations in Physics?

OpenStudy (anonymous):

I know the kinematics and -9.8 m/s^2 is gravity and that's about it

OpenStudy (anonymous):

the delta and stuff

OpenStudy (aum):

\(v^2 = u^2 - 2gx = 17^2 - 2*10*x = -20x + 289\) \(v = \sqrt{-20x + 289}\) where v is the velocity and x is the displacement (vertical height of the ball). I chose x because they want a plot of v vs displacement and so x for displacement will be convenient and v will go along the y-axis. Draw a graph of v vs x. When the ball reaches the maximum height, velocity v = 0. So the x-intercept of the graph will give the maximum height reached by the ball. PS: I am using g = 10 m/s^2